OpenStudy (pottersheep):

Please help me with grade 11 beginner trig :( Confused! Please help me with grade 11 beginner trig :( Confused! @Mathematics

OpenStudy (pottersheep):

Find 3 primary ratios for given angle. Delta = 135 degree

OpenStudy (saifoo.khan):

primary ratio = sin, cos tan?

OpenStudy (pottersheep):

yep

OpenStudy (saifoo.khan):

\[\sin 135 = \frac{1}{\sqrt2}\] is that the format u want?

OpenStudy (pottersheep):

Ohh, how did you get that? yes that's the format thank you!

OpenStudy (pottersheep):

do i do something on my calculator?

OpenStudy (saifoo.khan):

i did that on my calc. i knew the manual method as well which i forgot. LOL.

OpenStudy (pottersheep):

Oh that's okay then haha

OpenStudy (saifoo.khan):

\[\cos135 = -\frac{1}{\sqrt2}\]

OpenStudy (saifoo.khan):

do u want the working asw ell?

OpenStudy (saifoo.khan):

as well*

OpenStudy (pottersheep):

Yes pelase

OpenStudy (saifoo.khan):

\[\tan135 = -1\]

OpenStudy (saifoo.khan):

Lol, i dont knw the working.. sorry.

OpenStudy (pottersheep):

Then why did you ask LOL anyways thank you!

OpenStudy (saifoo.khan):

well, u can solve it by this way.

OpenStudy (cwrw238):